Missouri Journal of Mathematical Sciences Index

Author Index

Abotteen, Essam A., Inner Plethysm in the Representation Ring of the General Linear Group, Vol. 2, 115.

Abotteen, Essam, On the Structure of the Hopf Representation Ring of the Symmetry Groups, Vol. 4, 17.

Ahuja, Mangho, On Defining A Hyper-

bola, Vol. 1, No. 3, 23. Azarian, Mohammad K., On the Lower

Near Frattini Subgroups and Nearly Splitting Groups, Vol. 2, 18.

Azarian, Mohammad K., The Generating Function for the Fibonacci Sequence, Vol. 2, 78.

Azarian, Mohammad K., On the Lower Near Frattini Subgroups of Amalgamated Free Products of Groups I, Vol. 2, 105.

Azarian, Mohammad K., There May be More Than One Way to Find the Derivative of a Function, Vol. 5, 13.

Beach, David and Bobby Winters, Another Proof That the Closed Unit Interval is a Continuous Image of the Cantor Set, Vol. 5, 109.

Bennish, Joseph, An Application of the Unique Continuation Property to the Computation of a Fourier Transform, Vol. 4, 107.

Bennish, Joseph, On the Right–Hand Derivative of Integral Functions With a Discontinuous Integrand, Vol. 5, 4.

Bennish, Joseph, On the Reflection Properties of the Conic Sections, Vol. 5, 56.

Bennish, Joseph, Decomposition of the Line into Countably–Many Measure–Theoretic Dense Sets, Vol. 5, 123.

Bowman, Harold see Chou. Brown, Barry E. see Thuong, Vol. 1.

Cacioppo, Robert, Note on the First Fundamental Theorem for Riemann–Stieljtes Integrals, Vol. 4, 9.

Cammack, Larry see Schrag. Carter, Marte see Davenport. Chou, Wen–Seng, Harold Bowman, and

Jan–Shyong Shiue, A Matrix Method for Solving the Postage Stamps Problem, Vol. 3, 70.

Craven, Sherralyn, Some Highlights from ICME–6, Vol. 0, No. 1, 25. Davenport, Edward and Marte Carter,

Mathematics Placement and Assessment at Central Missouri State University, Vol. 4, 121.

Doucette, Robert L., A Challenging Area Problem Revisited, Vol. 5, 59. Edmondson, L. Vincent, Teaching About

Professionalism and Ethics in a Computer Science Course, Vol. 3, 61. Edmondson, L. Vincent, A Relationship Between the Metropolis Algorithm and the Two–Membered Evolution Strategy, Vol. 5, 8.

Eifler, Larry, A Simple Proof of Cauchy’s Integral Formula for Circles, Vol. 1, No. 2, 18.

Erickson, Martin, Shannon’s Theorem and the Binomial Random Variable, Vol 5, 2.

Euler, Russell, Evaluation of a Family of Improper Integrals, Vol. 0, No. 1, 3.

Euler, Russell, A Note on a Differential Equation, Vol. 1, No. 1, 26.

Euler, Russell, A Geometric Approach to a Trigonometric Integral, Vol. 1, No. 1, 28.

Euler, Russell, Fibonacci Decimal Number Patterns Via the Generating Function, Vol. 2, 137.

Volume 6, 1994 – Volume 10, 1998

Euler, Russell, Maximizing the Surface Area of N–Dimensional Unit Sphere, Vol. 4, 14.

Euler, Russell, Elementary Applications of

Complex Numbers, Vol. 4, 70. Euler, Russell, Evaluation of a Family of

Summations, Vol. 5, 17. Euler, Russell, A Generalized Exponential

Function, Vol. 5, 68. Foran, James and Paul Liebnitz, Expan-

sions of Baire Spaces, Vol. 1, No. 3, 2.

Francis, Richard L., Primes, Semi–Primes,

and Strong Composites, Vol. 2, 4. Francis, Richard L., Long–Ago Problems

for the Student of Today, Vol. 2, 51. Francis, Richard L., I Can’t See the Math

for the Problems, Vol. 3, 142. Francis, Richard L., Did You Do Anything

in Class Today?, Vol. 5, 21. Ganapathy, Jayanthi, Some Thoughts

on the Absolute Convergence of a Trigonometric Series, Vol. 3, 2.

Ganapathy, Jayanthi, An Example of a Function in Lip_{\frac{1}{2}} ([0, 2π]) With An Absolutely Non–Convergent Fourier Series, Vol. 5, 126.

Garimella, Ramesh, One–to–One Continuous Extensions of Analytic Functions, Vol. 2, 12.

Garimella, Ramesh, On the Common Zeroes of Finite Blaschke Products, Vol. 2, 70.

Garimella, Ramesh, A Counter Example

in Group Theory, Vol. 3, 77. Gunawardena, K. L. D., Stirling’s For-

mula: An Application of the Central Limit Theorem, Vol. 5, 107.

Hartzler, Stan, A Storming of our Pythagorean Aristocracy, Vol. 0, No. 1, 18.

Hicks, T. L., Extensions of Topologies,

Vol. 2, 62. John, David, Some Properties Related to Dendritic Spaces, Vol. 4, 116. John, David, A–Sets and Abcohesive

Spaces, Vol. 5, 63. Krantz, Steven G., On the Area Inside a

Circle, Vol. 4, 2. Larsen, Mogens E. and David Singmaster, Triangles with Equivalent Relations Between the Angles and Between the Sides, Vol. 3, 111. Liebnitz, Paul see Foran. Liu, Yi–Hsin and John Maloney, Discriminant Analysis Using A Bicriteria Linear Program, Vol. 4, 62. Liu, Yi–Hsin see Maloney, Vol. 4. Liu, Yi–Hsin see Maloney, Vol. 5. Livingston, William, Beginning Algebra:

Arithmetic in Disguise, Vol. 1, No. 1, 22.

Livingston, William, A Puzzle in Kelley’s

Appendix, Vol. 2, 26. Loper, K. Alan, On Unit Groups of Extension Rings: An Example, Vol. 3, 20.

Maloney, John and Yi–Hsin Liu, Using a Nonlinear Discriminant Function for Solving Discriminant Analysis Problems, Vol. 5, 83.

Maloney, John and Yi–Hsin Liu, A Quadratic Programming Approach for Solving Discriminant Analysis Problems, Vol. 4, 139. Maloney, John see Liu, Vol. 4. McDaniel, Wayne L., On the Intersection of the Sets of Base b Smith Numbers and Niven Numbers, Vol. 2, 132. Merryfield, Kent G., A Nowhere Analytic

C^{∞} Function, Vol. 4, 132. Moussavi, Saadat, A Generalization of Young’s Theorem, Vol. 4, 76.

Moussavi, Saadat, Accelerated MSOR

Method, Vol. 5, 24. Muthuvel, Kandasamy, Application of

Translation of Sets, Vol. 3, 139. Muthuvel, Kandasamy, A Note on Lower

Near Frattini Subgroups, Vol. 4, 11. Muthuvel, Kandasamy, A Note on a

Counter Example in Finite Groups, Vol. 5, 16.

Pirtle, Elbert, Noetherian Integrally Closed Duo Rings are Krull Rings, Vol. 1, No. 1, 9.

Rabinowitz, Stanley, The Volume of an n–Simplex With Many Equal Edges, Vol. 1, No. 2, 11.

Rabinowitz, Stanley, A Model for Memory Interference in Multiprocessors, Vol. 3, 12.

Redmond, Don, Consecutive Composite Values of a Quadratic Polynomial, Vol. 3, 25.

Rhee, Noah H., Calculating An Invariant Subspace of Diagonally Dominant Matrices – Part I, Vol. 5, 71.

Rhee, Noah H., Calculating An Invariant Subspace of Diagonally Dominant Matrices – Part II, Vol. 5, 116.

Schrag, Gerald and Larry Cammack, On the t–Extendability of the Generalized Petersen Graphs, Vol. 1, No. 1, 3.

Shiue, Jan–Shyong see Chou. Singmaster, David see Larsen. Skow, Donald P. see Wiener, Vol. 3. Skow, Donald P. see Wiener, Vol. 4. So, Shing–Seung, Almost Locally Con-

nected (SO) Spaces, Vol. 0, No. 1, 7.

So, Shing–Seung, Convergence Semi-

groups, Vol. 1, No. 3, 11. So, Shing–Seung, Sequences, Mathemat-

ical Induction, and the Heine–Borel Theorem, Vol. 3, 84.

So, Shing S. see Tsang. Thuong, Tran van, Homotopy Exten-

sion Theorem for a Fiber–Preserving Piecewise Linear μ–Homotopy, Vol. 3, 33.

Thuong, Tran van, Some Applications of the Stone–Weierstrass Theorem, Vol. 3, 79.

Thuong, Tran van and Barry E. Brown, How Mathematics Grows, Vol. 1, No. 3, 28.

Travis, Betty, Breaking the “Zip Code”

Code, Vol. 1, No. 2, 2. Tsang, Philip M., and Shing S. So, Gen-

erating Functions: Applications and Techniques, Vol. 4, 109.

Vayo, H. Westcott, Model Geometry For

HIV–1 and –2, Vol. 4, 49. Walters, F. Garnett, Computer Science is

Alive and Well on Planet Earth, Vol. 1, No. 2, 23.

Watkins, Will see Wiener, Vol. 4. Wiener, Bella and Joseph Wiener, The

Analytical Machinery of Symmetry, Vol. 3, 130.

Wiener, Bella and Joseph Wiener, A Trigonometric Derivation of Point–To–Line and Point–To–Plane Distance Formulas, Vol. 5, 113.

Wiener, Joseph, An Analytical Approach to a Trigonometric Integral, Vol. 2, 75.

Wiener, Joseph, Resonance in Linear Differential Equations and L’Hospital’s Rule, Vol. 2, 96.

Wiener, Joseph, see Wiener, Bella, Vol. 3. Wiener, Joseph see Wiener, Bella, Vol. 5. Wiener, Joseph and Donald P. Skow, Dif-

ferentiating Indefinite Integrals with Respect to a Parameter, Vol. 3, 65.

Wiener, Joseph, Donald Skow, and Will Watkins, Integrating Powers of

Trigonometric Functions, Vol. 4, 55. Winters, Bobby see Beach. Abotteen, Essam A., Inner Plethysm in

the Representation Ring of the General Linear Group, Vol. 2, 115.

Abotteen, Essam, On the Structure of the Hopf Representation Ring of the Symmetry Groups, Vol. 4, 17.

Ahuja, Mangho, On Defining A Hyper-

bola, Vol. 1, No. 3, 23. Azarian, Mohammad K., On the Lower

Near Frattini Subgroups and Nearly Splitting Groups, Vol. 2, 18.

Azarian, Mohammad K., The Generating Function for the Fibonacci Sequence, Vol. 2, 78.

Azarian, Mohammad K., On the Lower Near Frattini Subgroups of Amalgamated Free Products of Groups I, Vol. 2, 105.

Azarian, Mohammad K., There May be More Than One Way to Find the Derivative of a Function, Vol. 5, 13.

Beach, David and Bobby Winters, Another Proof That the Closed Unit Interval is a Continuous Image of the Cantor Set, Vol. 5, 109.

Bennish, Joseph, An Application of the Unique Continuation Property to the Computation of a Fourier Transform, Vol. 4, 107.

Bennish, Joseph, On the Right–Hand Derivative of Integral Functions With a Discontinuous Integrand, Vol. 5, 4.

Bennish, Joseph, On the Reflection Properties of the Conic Sections, Vol. 5, 56.

Bennish, Joseph, Decomposition of the Line into Countably–Many Measure–Theoretic Dense Sets, Vol. 5, 123.

Bowman, Harold see Chou.

Brown, Barry E. see Thuong, Vol. 1.

Cacioppo, Robert, Note on the First Fundamental Theorem for Riemann–Stieljtes Integrals, Vol. 4, 9.

Cammack, Larry see Schrag. Carter, Marte see Davenport. Chou, Wen–Seng, Harold Bowman, and

Jan–Shyong Shiue, A Matrix Method for Solving the Postage Stamps Problem, Vol. 3, 70.

Craven, Sherralyn, Some Highlights from ICME–6, Vol. 0, No. 1, 25. Davenport, Edward and Marte Carter,

Mathematics Placement and Assessment at Central Missouri State University, Vol. 4, 121.

Doucette, Robert L., A Challenging Area Problem Revisited, Vol. 5, 59. Edmondson, L. Vincent, Teaching About

Professionalism and Ethics in a Computer Science Course, Vol. 3, 61. Edmondson, L. Vincent, A Relationship Between the Metropolis Algorithm and the Two–Membered Evolution Strategy, Vol. 5, 8.

Eifler, Larry, A Simple Proof of Cauchy’s Integral Formula for Circles, Vol. 1, No. 2, 18.

Erickson, Martin, Shannon’s Theorem and the Binomial Random Variable, Vol 5, 2.

Euler, Russell, Evaluation of a Family of Improper Integrals, Vol. 0, No. 1, 3.

Euler, Russell, A Note on a Differential Equation, Vol. 1, No. 1, 26.

Euler, Russell, A Geometric Approach to a Trigonometric Integral, Vol. 1, No. 1, 28.

Euler, Russell, Fibonacci Decimal Number Patterns Via the Generating Function, Vol. 2, 137.